Evaluate 4x^2+3 when x=5
Evaluate 2a-3 when x=-5
Evaluate -2a+6b when a=6 and b=-3

Answer by Alan3354(69443)

Evaluate 4x^2+3 when x=5
Sub 5 for x
= 4*5^2+3
= 4*25+3
= 103
===================
The others are similar.
---
Evaluate 2a-3 when x=-5
Evaluate -2a+6b when a=6 and b=-3